OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of -1 and passes through the point (-2, 5)

OpenStudy (anonymous):

can somone remind me whta the point slope formula is again?

OpenStudy (anonymous):

y = mx + b

OpenStudy (anonymous):

no - thats slope intercept point slope is y-y1 = m (x-x1) i'm from uk and we dont usually use these terms

OpenStudy (anonymous):

its all good

OpenStudy (anonymous):

so from your info y - 5 = -1(x + 1)